OpenStudy (queenbb):

I Will give medal 1. Delaney would like to make a 5 lb nut mixture that is 60% peanuts and 40% almonds. She has several pounds of peanuts and several pounds of a mixture that is 20% peanuts and 80% almonds. Let p represent the number of pounds of peanuts needed to make the new mixture, and let m represent the number of pounds of the 80% almond-20% peanut mixture. (a) What is the system that models this situation? (b) Which of the following is a solution to the system: 2 lb peanuts and 3 lb mixture; 2.5 lb peanuts and 2.5 lb mixture; 4 lb peanuts and 1 lb mixture? Show your work. Answer (Score for Question 2: ___ of 5 points) 2. Seiji and Gavin both worked hard over the summer. Together they earned a total of $425. Gavin earned $25 more than Seiji. (a) Write a system of equations for the situation. Use s for the amount Seiji earned and g for the amount Gavin earned. (b) Graph the equations in the system. (c) Use your graph to estimate how much each person earned. Answer: (Score for Question 3: ___ of 5 points) 3. Two sidewalks in a park are represented by lines on a coordinate grid. Two points on each of the lines are shown in the tables. Sidewalk 1 x y 2 7 0 3 Sidewalk 2 x y 1 5 3 3 (a) Write the equation for Sidewalk 1 in slope-intercept form. (b) Write the equation for Sidewalk 2 in point-slope form and then in slope-intercept form. (c) Is the system of equations consistent independent, coincident, or inconsistent? Explain. (d) If the two sidewalks intersect, what are the coordinates of the point of intersection? Use the substitution method and show your work.

OpenStudy (queenbb):

can you guys help

OpenStudy (queenbb):

I need help

OpenStudy (queenbb):

hello plz help me me

OpenStudy (queenbb):

@welshfella

OpenStudy (queenbb):

help

OpenStudy (welshfella):

First of all p + m = 5 because there is a total of 5 pounds of mixture to be made. Now for there to be 40% of almonds there must be 0.4*5 = 2 pounds almonds So as there are 80% almonds in the mixture the amount of m must be 2 / 0.8 = 2.5 pounds so we have p + 2.5 = 5 so p = 2.5 Answer is 2.5 pounds of peanuts and 2.5 pounds of the mixture.

OpenStudy (queenbb):

which answer is that one for

OpenStudy (welshfella):

question 1

OpenStudy (queenbb):

ok what about question two an three

OpenStudy (queenbb):

@welshfella

OpenStudy (welshfella):

i goota go sorry

OpenStudy (queenbb):

ok

OpenStudy (queenbb):

@AloneS

alones (alones):

Make a new post for number \(2\)

OpenStudy (queenbb):

ok